Thomas’ Calculus Early Transcendentals 13ed

Homework Assignment for Chapter 04

1. Section 4.1: problems 39 (need not graph it), 41, 49, 67 (absolute extreme values only), 71 (absolute extreme values only), 77.

2. Section 4.2: problems 18, 19, 21 (see Example 1), 31, 39, 61, 65.

3. Section 4.3: problems 7, 11, 27, 33, 43, 51, 74, 77.

4. Section 4.4: problems 77, 85, 105, 111.

5. Section 4.5: problems 21, 39, 41, 51, 57, 61, 63, 65, 69, 71, 73, 80, 81(b), 84(c), 86(c,d), 88. Remark: In problem 69, 71, 73, find the correct limit instead.

6. Section 4.6: problems 12, 27, 67(a).

7. Section 4.8: Problems 79, 80, 84, 95, 103, 107, 109 127, 128.